OpenStudy (anonymous):

estimate the quantity using linear approximation 1/sqrt (99)- 1/sqrt(98)

OpenStudy (shayaan_mustafa):

Hi andrea723 :) are you there sir?

OpenStudy (anonymous):

here

OpenStudy (shayaan_mustafa):

\[\frac{1}{\sqrt{99}}-\frac{1}{\sqrt{98}}\] Is this your question look like?

OpenStudy (anonymous):

yes

OpenStudy (shayaan_mustafa):

do you know the formula for linear approximation?

OpenStudy (anonymous):

yes f(a)+f^1(a)(x-a)?

OpenStudy (shayaan_mustafa):

hello :) are yo still there?

OpenStudy (anonymous):

yes sorry

OpenStudy (shayaan_mustafa):

formula for linear approximation is given as.\[\large f(x)\approx f(x_0)-f'(x_0)(x-x_0)\]

OpenStudy (shayaan_mustafa):

find derivative of your function. and tell me.

OpenStudy (anonymous):

-1/2x^(-3/2)?

OpenStudy (shayaan_mustafa):

OK. what did you choose x? and also a?

OpenStudy (anonymous):

isnt it 99 and 98?

OpenStudy (shayaan_mustafa):

yes x=99 and 98 what about a?

OpenStudy (anonymous):

100??

OpenStudy (shayaan_mustafa):

Hi andrea723 first we have to calculate separate for 1/sqrt(99) and then 1/sqrt(98) and then we will subtract both answers. That will be your require answer.

OpenStudy (anonymous):

i got -5.11e^(-4)?

OpenStudy (shayaan_mustafa):

is it all your answer?

OpenStudy (shayaan_mustafa):

If so then your answer is right.

OpenStudy (anonymous):

i dont know when i submit that answer online is doesnt work thank you so much anyway

OpenStudy (shayaan_mustafa):

If you still have confusion then you can ask me. But I am pretty sure that your answer is correct even you can ask it anyone here. Good luck friend :)